# Manipulating the Online Marketplace of Ideas

### v4

@@ Abstract
Social media, seen by some as the modern public square, is vulnerable to manipulation. By controlling inauthentic accounts impersonating humans, malicious actors can amplify disinformation within target communities. The consequences of such operations are difficult to evaluate due to the challenges posed by collecting data and carrying out ethical experiments that would influence online communities. Here we use a social media model that simulates information diffusion in an empirical network to quantify the impacts of several adversarial manipulation tactics on the quality of content. We find that the presence of influential accounts, a hallmark of social media, exacerbates the vulnerabilities of online communities to manipulation. Among the explored tactics that bad actors can employ, infiltrating a community is the most likely to make low-quality content go viral. Such harm can be further compounded by inauthentic agents flooding the network with low-quality, yet appealing content, but is mitigated when bad actors focus on specific targets, such as influential or vulnerable individuals. These insights suggest countermeasures that platforms could employ to increase the resilience of social media users to manipulation.
Significance Statement: We show that social media users are vulnerable to adversarial manipulation tactics, through which bad actors can amplify exposure to content that potentially threatens democratic elections and public health. While tactics such as flooding the network with low-quality yet appealing content are damaging, getting users to follow inauthentic accounts has the most detrimental impact. Bad actors can increase harm by maximizing coverage rather than targeting particular individuals, such as influential ones. The varying degrees of harm associated with these tactics highlight tradeoffs and specific areas on which platforms could focus as they develop deterrence strategies against manipulation.
@@ 1 Introduction
The vision of social media as the modern public square has been challenged as users have become victims of manipulation by astroturf (1, 2), trolling (3), impersonation (4), and misinformation (5, 6, 7). False news have been reported to spread virally — similarly to reliable information (8) or even more (9) depending on operational definitions. These kinds of manipulation exploit a complex interplay of socio-cognitive (10, 11), ideological (7), and algorithmic (12, 13) biases. The exploitation is enabled or greatly facilitated by inauthentic accounts that impersonate people with malicious intent. Many such accounts can be coordinated by a single entity (14), either manually or through software applications commonly known as social bots (15, 16). Inauthentic and/or coordinated accounts have been observed to amplify disinformation (8), influence public opinion (17, 18, 19, 14), commit financial fraud (20, 14), infiltrate vulnerable communities (21, 3, 22), and disrupt communication (23, 24).
Can social media be manipulated to the point that they no longer function as a public square? Under what conditions? It is difficult to carry out empirical experiments and analyses in the real world to explore these questions (25, 26). One challenge is the limited size of experiments in the wild, stemming from both costs and ethical concerns about the potentially harmful nature of content from bad actors. A second difficulty is the limited data from social media platforms available to researchers (27), exacerbated by recent events such as the acquisition of Twitter by Elon Musk. These difficulties have led, for example, to conflicting accounts about whether disinformation campaigns on social media can sway elections (28, 29, 30, 31, 32, 33). Evidence suggests that these operations mainly impact specific vulnerable communities (34, 35). However, we lack a comprehensive quantitative understanding of how coordinated inauthentic tactics can disrupt online communities. This prevents the informed design of moderation or regulatory policies to protect the online public square from adversarial manipulation.
@@
We find that the presence of manipulation is sufficient to suppress quality information, driving low-quality content to spread virally in the network. We also examine network vulnerabilities that may amplify the effects of manipulation, and evaluate the overall information quality as the result of different malicious tactics, such as infiltrating a community, generating attention-grabbing content, flooding the network, and targeting specific individuals. Insights from these analyses are instrumental in developing countermeasures to increase the resilience of social media and their users against manipulation. We discuss mitigation steps that platforms could take and the issues that arise from regulations aimed at protecting human speech from suppression.
@@ Results
Even though people prefer quality content (36), their sharing behavior is mediated by other factors such as laziness (37) and message appeal. To account for this, each message $m$ in the model has two intrinsic and independent attributes. The appeal $a_{m}$ models the likelihood that the message is actually reshared by agents (Fig. 1). The quality $q_{m}$, on the other hand, represents objective, desirable properties of content such as the originality of an idea or the accuracy of a claim. Here we naively represent quality as a scalar value. Deceptive posts may have low quality yet high appeal. For example, false news and junk science articles have low quality — most people would not share them knowingly. Yet such low-quality content may be even more likely to spread virally than high-quality information (9). Low-quality content may be novel, clickbait, ripped from headlines, and/or may appeal to people’s political, emotional, or conspiratorial bias. Worse yet, bad actors can employ generative AI to produce such content at scale (38).
@@
Infiltration describes how bad actors amplify exposure to their messages by getting authentic accounts to follow them (Fig. 2(a)). Bad actor infiltration into the social network is modeled by a parameter $\gamma$, the probability that a bad actor is followed by an authentic agent. Unless otherwise stated, we assume that authentic agents follow bad actors uniformly at random. Fig. 2(b) illustrates the effective suppression of information quality when $\gamma$ is high.
@@
Figure 2: Subnetworks modeling authentic accounts (purple nodes) and bad actors (yellow nodes). (a) Illustration of the follower link structure. Solid links indicate follower relations within each subnetwork. Both subnetworks have hub and clustering structure that mimics or derives from online social networks. Dashed links represent authentic accounts following bad actors, according to the infiltration parameter $\gamma$, which represents the probability that an authentic node follows any given bad actor. When $\gamma=0$ there is no infiltration and bad actors are isolated, therefore harmless; the opposite extreme $\gamma=1$ indicates complete infiltration, such that bad actors dominate the network. (b) Effects of bad actor infiltration $\gamma$ on the quality of messages in synthetic networks with $10^{3}$ authentic agents and 100 inauthentic agents. For illustration purposes, both the authentic and inauthentic subnetworks in this panel are generated with the same method used for the inauthentic subnetworks in our experiments (see Methods). Node size represents the number of followers. The darker an authentic agent node, the lower the quality of messages in their feed.
@@
The quality $q$ and appeal $a$ of messages originating from authentic accounts are drawn independently from two distinct distributions, reflecting empirical evidence that these messages tend to have high quality and low appeal (see Supplementary Material). In contrast, we assume that bad actors can manipulate information in the network by creating messages with low quality ($q=0$) and deceptively high appeal. The appeal differential of content from bad actors is modeled by the deception parameter $\phi$, defined as the probability that bad actor content is irresistible ($a=1$). In the absence of deception ($\phi=0$), the appeal of bad actor messages is drawn from the same distribution as those from authentic accounts (see details in Methods).
@@
Flooding is another tactic inauthentic accounts can use to amplify their influence, by crowding out high-quality information. To model this, the parameter $\theta$ is defined as the ratio between the exposure of bad actor content and authentic content (see Methods). Bad actors can achieve high exposure in several ways, including posting at high frequency/volume and artificially inflating popularity/engagement indicators (8, 40, 38).
@@
SimSoM lets us explore information diffusion on social media, including the properties of authentic accounts that might render them vulnerable to adversarial attacks and the effects of different manipulation tactics. In the next sections, we present the results from simulations of the model on online communities derived from an empirical follower network ($N\approx 10^{4}$ Twitter accounts). The network has both scale-free structure (hubs) and high clustering (triangles), structural characteristics that are ubiquitous in socio-technical networks. It also has a realistic community structure, with two well-separated groups of accounts capturing political polarization (see Methods).
@@
Once the system reaches a steady state, in which the message quality across the network has stabilized, we record the mean quality of the messages in the feeds of authentic agents. These measurements are further averaged across simulation runs with the same parameters but different random seeds. We simulate the information diffusion process in networks with different structures to explore whether social network features render communities vulnerable to bad actor tactics. Similarly, we evaluate the impact of these tactics by evaluating the model with varying levels of bad actor infiltration ($\gamma$), deception ($\phi$), flooding ($\theta$), and targeting specific types of accounts. We report on the relative quality, defined as the ratio of the average quality of authentic agents to that of a baseline without bad actors (see Methods).
@@
Figure 3: Impacts of different network structural features on the average information quality, relative to the scenario without bad actors. The original network (“hubs + communities”) is visualized along with shuffled networks in which links from the original network are rewired while preserving communities, hubs, or neither (“random”). Node size and color represent, respectively, the number of followers of an account and their political leaning ranging from liberal to conservative (red to blue, see Methods). Yellow nodes are bad actors. Pairwise statistical significance is calculated using Welch’s two-sided t-test (**** for $p<10^{-4}$); only significant differences are reported.
Key structural features of the social network may play a role in amplifying our vulnerability to manipulation by bad actors. The empirical network has two features that are ubiquitous in social media: the presence of hubs and highly clustered communities. We can explore how overall quality is affected by these features through three alternative networks constructed by shuffling links while preserving hubs, community structure, or neither (see Methods). Fig. 3 shows that community structure does not significantly affect the overall quality (purple vs. blue) because we assume that agents in each community are equally likely to follow bad actors. On the other hand, having hubs makes a network more vulnerable to manipulation: the relative system quality is significantly lower in the presence of hubs, both when there are communities (purple vs. green, $p<10^{-4}$) and not (blue vs. orange, $p<10^{-4}$). This is because node in-degrees and out-degrees are highly correlated (Spearman correlation 0.9, $p<10^{-4}$) in the empirical network. Therefore, when authentic followers are concentrated among hubs, high-quality content is also concentrated among those hubs. This implies that high-quality content has more competition and becomes obsolete more quickly compared to the case in which this content is uniformly distributed among authentic nodes. The same does not apply to content from bad actors because this content spreads uniformly among authentic users. (The scenario in which content from bad actors mostly concentrates among hubs is explored later.)
@@
Bad actors may maximize their message spread by combining various manipulation tactics. We systematically quantify the effects of these tactics through simulations varying the parameters for infiltration ($10^{-4}\leq\gamma\leq 10^{-1}$), deception ($0\leq\phi\leq 1$), and flooding ($1\leq\theta\leq 32$). See Methods for further details. Fig. 4(a,b,c) illustrates the effects of individual malicious tactics on the overall quality of information spreading through the network, compared to a baseline scenario without bad actors. We observe that infiltration is the most harmful manipulation tactic: when authentic agents have a $\gamma=10\%$ probability of following each bad actor, the average quality in the system is reduced to less than half. Flooding and deception have smaller effects. When low-quality content has $\theta=32$ times more exposure than authentic content, quality is reduced to less than 70%. When bad actors generate content with maximum appeal exclusively ($\phi=1$), quality is reduced to about 70%.
@@
Similarly, Fig. 4(d,e,f) shows the effects of pairs of tactics combined. Infiltration is dominant, but more harm can be done in combination with flooding or deception: the average quality is reduced to 40% when $\gamma=0.1$ and $\theta=32$ (Fig. 4(d)) or $\phi=1$ (Fig. 4(e)). Combining flooding and deception ($\theta=32,\phi=1$, Fig. 4(f)) only results in marginal loss of quality (below 70%). With all three tactics combined ($\gamma=0.1,\theta=32,\phi=1$, not shown), bad actors can further reduce the quality to 30%.
@@
Empirical evidence has shown that among fact-checked claims, low-quality content (debunked claims) tends to have larger retweet cascades than high-quality content (confirmed claims) (9, 41). SimSoM allows us to examine how different factors may contribute to such a virality pattern. Fig. 5 illustrates the effects of bad actor tactics on the size of reshare cascades for content generated by both authentic agents (“high-quality”) and inauthentic agents (“low-quality”). We observe that low-quality content can be boosted most effectively through high bad actor infiltration. This appears to be at odds with the hypothesis that attributes the empirical difference in virality to factors like novelty, which make false news more appealing, rather than to inauthentic accounts such as social bots (9). In our model, even the highest boost in appeal ($\phi=1$) does not grant low-quality cascades a virality comparable to those under higher bad actor amplification. A couple of observations reconcile the apparent contradiction. First, multiple factors may contribute to the virality of low-quality information, including inauthentic accounts and deceptively appealing content. Second, data from Twitter does not allow for the reconstruction of actual cascades. Therefore, empirical analyses may underestimate intermediary amplification by inauthentic accounts even when those accounts are removed. On the other hand, our model lets us reconstruct the likely reshare cascades that include intermediary amplification by inauthentic accounts.
Figure 5: Complementary cumulative distributions of reshare cascade sizes for low- and high-quality content, generated by inauthentic and authentic agents, respectively. The plots are based on single runs of the model. (a) Effect of bad actor infiltration $\gamma$, with no flooding ($\theta=1$) or deception ($\phi=0$). (b) Effect of flooding $\theta$, with low infiltration ($\gamma=10^{-3}$) and no deception ($\phi=0$). (c) Effect of deception $\phi$, with low infiltration ($\gamma=10^{-3}$) and no flooding ($\theta=1$).
To date, empirical data from social media platforms has allowed researchers to measure reshare cascades, but not exposure. Using the SimSoM model, it is possible to reconstruct not only likely reshare networks but also exposure networks, thus estimating how many accounts are exposed to (i.e., view) a message even if they do not reshare it (see Methods). Fig. 6 compares the sizes of reshare and exposure cascades. In general, reshares underestimate exposures by roughly one order of magnitude. Excluding the smallest and largest cascades, we observe that exposure networks grow sub-linearly with reshare networks: $s_{v}\sim s_{r}^{\nu}$ where $s_{v}$ and $s_{r}$ are the exposure and reshare cascade sizes, respectively, and $\nu<1$ is the scaling exponent. This means that as messages go viral, exposures grow more slowly than reshares. The exponent is higher for low-quality ($\nu\approx 0.8$) than for high-quality content ($\nu\approx 0.6$), suggesting that for each extra reshare, messages posted by inauthentic agents gain more views.
@@
As an example, an adversary might target influentials based on the assumption that having such followers can multiply their impact — a message reshared by an influential account has a higher chance of going viral. Targeting influentials is well within the capability of bad actors and even automated accounts; the number of followers, often used as a proxy for influence (42), is public information on all social media platforms. A bad actor can easily interact with accounts having many followers by mentioning and/or following them (43, 44); other ploys include retweeting, quoting, and/or liking their tweets. There is empirical evidence of preferential targeting by bad actors that spread misinformation (1, 8).
Targeting politically active accounts or habitual misinformation spreaders are also conceivable tactics. An important question, then, is whether tactics targeting specific authentic accounts do in fact increase the manipulative power of bad actors. To explore this question, we introduce a preferential targeting tactic for adversarial actors in the model. Authentic agents have different probabilities of following inauthentic ones, proportional to one of five features that are available in the empirical data: number of followers (hubs tactic), propensity to share misinformation (misinformation tactic), political partisanship (partisanship tactic), or specific political leaning (liberal and conservative tactic). See Methods for details.
Figure 7: Effects of targeting tactics. (a) Average information quality resulting from each tactic, as well as the default random targeting, relative to the scenario without bad actors. We highlight significant differences calculated using Welch’s two-sided t-test (*** for $p<10^{-3}$ and **** for $p<10^{-4}$). (b) Suppression of quality in the empirical network when bad actors specifically target influential accounts (hubs), and when they target politically left- (liberal) and right-leaning (conservative) accounts. The network has $10^{3}$ authentic agents (purple nodes) and 50 inauthentic agents. Node size represents the number of followers. The darker an authentic agent node, the lower the quality of messages in their feed. Significant changes due to targeting tactics are only observed when bad actor infiltration is sufficiently high, therefore we use $\gamma=10^{-1}$ in experiments for both panels.
@@
The harm of manipulation through the network also diminishes when bad actors target accounts sharing a lot of misinformation ($p<10^{-3}$) or with specific political leaning ($p<10^{-4}$). The echo-chamber structure of the empirical network (Fig. 7(b)) helps interpret the latter finding: low-quality messages get shared and become obsolete rapidly within one densely connected partisan sub-community, sparing the rest of the network.

### v5

# Quantifying the Vulnerabilities of the Online Public Square to Adversarial Manipulation Tactics

@@ Abstract
Social media, seen by some as the modern public square, is vulnerable to manipulation. By controlling inauthentic accounts impersonating humans, malicious actors can amplify disinformation within target communities. The consequences of such operations are difficult to evaluate due to the challenges posed by collecting data and carrying out ethical experiments that would influence online communities. Here we use a social media model that simulates information diffusion in an empirical network to quantify the impacts of adversarial manipulation tactics on the quality of content. We find that the presence of hub accounts, a hallmark of social media, exacerbates the vulnerabilities of online communities to manipulation. Among the explored tactics that bad actors can employ, infiltrating a community is the most likely to make low-quality content go viral. Such harm can be further compounded by inauthentic agents flooding the network with low-quality, yet appealing content, but is mitigated when bad actors focus on specific targets, such as influential or vulnerable individuals. These insights suggest countermeasures that platforms could employ to increase the resilience of social media users to manipulation.
Significance Statement: We show that social media users are vulnerable to adversarial manipulation tactics, through which bad actors can amplify exposure to content that threatens, for example, democratic elections and public health. While tactics such as flooding the network with low-quality yet appealing content are damaging, getting users to follow inauthentic accounts has the most detrimental impact. Bad actors can increase harm by maximizing coverage rather than targeting particular individuals, such as influential ones. The varying degrees of harm associated with these tactics highlight tradeoffs and specific areas on which platforms could focus as they develop deterrence strategies against manipulation.
**footnotetext: To whom correspondence should be addressed: baotruon@iu.edu
@@ 1 Introduction
The vision of social media as the modern public square has been challenged as users have become victims of manipulation by astroturf (1, 2), trolling (3), impersonation (4), and misinformation (5, 6, 7). False news have been reported to spread virally — similarly to reliable information (8) or even more (9) depending on operational definitions. These kinds of manipulation exploit a complex interplay of socio-cognitive (10, 11), ideological (7), and algorithmic (12, 13) biases. The exploitation is enabled or greatly facilitated by inauthentic accounts that impersonate people with malicious intent. Many such accounts can be coordinated by a single entity (14). If this is done through software, such accounts are commonly referred to as social bots (15, 16). Inauthentic and/or coordinated accounts have been observed to amplify disinformation (8), influence public opinion (17, 18, 19, 14), commit financial fraud (20, 14), infiltrate vulnerable communities (21, 3, 22), and disrupt communication (23, 24).
Can social media be manipulated to the point that they no longer function as a public square? Under what conditions? It is difficult to carry out empirical experiments and analyses in the real world to explore these questions (25, 26). One challenge is the limited size of experiments in the wild, stemming from both costs and ethical concerns about the potentially harmful nature of content from bad actors. A second difficulty is the limited data from social media platforms available to researchers (27), exacerbated by recent events such as changes at Twitter/X. These difficulties have led, for example, to conflicting accounts about whether disinformation campaigns on social media can sway elections (28, 29, 30, 31, 32, 33). Evidence suggests that these operations mainly impact specific vulnerable communities (34, 35). However, we lack a comprehensive quantitative understanding of how coordinated inauthentic tactics can disrupt online communities. This prevents the informed design of moderation or regulatory policies to protect the online public square from manipulation.
@@
We find that the presence of manipulation is sufficient to suppress quality information, driving low-quality content to spread virally in the network. We also examine network vulnerabilities that may amplify the effects of manipulation, and evaluate the overall information quality as a result of different malicious tactics. We focus on four well-documented tactics commonly employed in influence operations across various platforms (36): (i) infiltrating a community, for example through social bots (6, 37), follow trains (38), or by impersonating news outlets (36, 37); (ii) generating deceptively appealing content, such as novel narratives (9) or emotional messages (39); (iii) flooding the network with high volumes of content by posting at high frequency to artificially inflate popularity/engagement indicators (8, 37, 40), and possibly deleting content to avoid detection (41); and (iv) targeting specific users, such as influential (8) or vulnerable individuals (6). Insights from analyses of these tactics are instrumental in developing countermeasures to increase the resilience of social media and their users against manipulation. We discuss mitigation steps that platforms could take and the issues that arise from regulations aimed at protecting human speech from suppression.
@@ Results
Even though people prefer quality content (42), their sharing behavior is mediated by other factors such as laziness (43) and message appeal. To account for this, each message $m$ in the model has two intrinsic and independent attributes, appeal $a_{m}$ and quality $q_{m}$. Everything else being equal, messages with higher appeal are more likely to be reshared (Fig. 1). Quality, on the other hand, represents objective, desirable properties of content such as the originality of an idea or the accuracy of a claim. Here we naively represent quality and appeal as scalar values. Deceptive posts may have low quality yet high appeal. For example, false news and junk science articles have low quality — most people would not share them knowingly. Yet such low-quality content may be even more likely to spread virally than high-quality information (9). Low-quality content may be novel, clickbait, ripped from headlines, and/or may appeal to people’s political, emotional, or conspiratorial bias. Worse yet, bad actors can employ generative AI to produce such content at scale (37).
@@
Infiltration describes how bad actors amplify exposure to their messages by getting authentic accounts to follow them (Fig. 2(a)). Bad-actor infiltration of the social network is modeled by a parameter $\gamma$, the probability that each bad actor is followed by each authentic agent. Unless otherwise stated, we assume that authentic agents follow bad actors uniformly at random. Fig. 2(b) illustrates the effective suppression of information quality when $\gamma$ is high.
@@
Figure 2: Subnetworks modeling authentic accounts (purple nodes) and bad actors (yellow nodes). (a) Illustration of the follower link structure. Solid links indicate follower relations within each subnetwork. Both subnetworks have hub and clustering structure that mimics or derives from online social networks. Dashed links represent authentic accounts following bad actors, according to the infiltration parameter $\gamma$, which represents the probability that an authentic node follows any given bad actor. When $\gamma=0$ there is no infiltration and bad actors are isolated, therefore harmless; the opposite extreme $\gamma=1$ indicates complete infiltration, such that bad actors are followed by all authentic accounts. (b) Effects of bad-actor infiltration $\gamma$ on the quality of messages in synthetic networks with $10^{3}$ authentic agents and 100 inauthentic agents. For illustration purposes, both the authentic and inauthentic subnetworks in this panel are generated with the same method used for the inauthentic subnetworks in our experiments (see Methods). Node size represents the number of followers. The darker an authentic agent node, the lower the quality of messages in their feed.
@@
The quality $q$ and appeal $a$ of messages originating from authentic accounts are drawn independently from two distinct distributions, reflecting empirical evidence that these messages tend to have high quality and low appeal (see Supplementary Material). In contrast, we assume that bad actors can manipulate information in the network by creating messages with low quality ($q=0$) and deceptively high appeal. The appeal differential of content from bad actors is modeled by the deception parameter $\phi$. In the absence of deception ($\phi=0$), the appeal of bad-actor messages is drawn from the same distribution as those from authentic accounts. If $\phi>0$, it represents the probability that bad-actor content has $a=1$, making it irresistible (see details in Methods).
@@
Flooding is another tactic inauthentic accounts can use to amplify their influence, by crowding out high-quality information. To model this, the parameter $\theta$ is defined as a boost of engagement for bad-actor content (see Methods).
@@
SimSoM lets us explore information diffusion on social media, including the properties of authentic accounts that might render them vulnerable to adversarial attacks and the effects of different manipulation tactics. In the next sections, we present the results from simulations of the model on online communities derived from an empirical follower network ($N\approx 10^{4}$ Twitter accounts). The network has both scale-free structure (hubs) and high clustering (triangles), structural characteristics that are ubiquitous in socio-technical networks. It also has a realistic community structure, with two well-separated clusters of accounts capturing political polarization (see Methods).
@@
Once the system reaches a steady state, in which the message quality across the network has stabilized (see Methods), we record the mean quality of the messages in the feeds of authentic agents. These measurements are further averaged across simulation runs with the same parameters but different random seeds. We simulate the information diffusion process in networks with different structures to explore whether social network features render communities vulnerable to bad-actor tactics. Similarly, we evaluate the impact of these tactics by evaluating the model with varying levels of bad-actor infiltration ($\gamma$), deception ($\phi$), flooding ($\theta$), and targeting specific types of accounts. We report on the relative quality, defined as the ratio of the average quality of authentic agents to that of a baseline without bad actors (see Methods).
@@
Figure 3: Impacts of different network structural features on the average information quality, relative to the scenario without bad actors. The original network (“hubs + clusters”) is visualized along with shuffled networks in which links from the original network are rewired while preserving clusters, hubs, or neither (“random”). Node size and color represent, respectively, the number of followers of an account and their political leaning ranging from liberal to conservative (red to blue, see Methods). Yellow nodes are bad actors. Pairwise statistical significance is calculated using the Mann–Whitney U test (*** for $p<10^{-3}$); only significant differences are reported.
Key structural features of the social network may play a role in amplifying our vulnerability to manipulation by bad actors. The empirical network has two features that are ubiquitous in social media: the presence of hubs and highly clustered communities. We can explore how overall quality is affected by these features through three alternative networks constructed by shuffling links while preserving hubs, cluster structure, or neither (see Methods). Fig. 3 shows that the presence of clusters does not significantly affect the overall quality (purple vs. blue) because we assume that agents in each cluster are equally likely to follow bad actors. On the other hand, having hubs makes a network more vulnerable to manipulation: the relative system quality is significantly lower in the presence of hubs, both when there are clusters (purple vs. green, $p<10^{-3}$) and not (blue vs. orange, $p<10^{-3}$). This is because node in-degrees and out-degrees are highly correlated (Spearman correlation 0.9, $p<10^{-4}$) in the empirical network. Therefore, when authentic followers are concentrated among hubs, high-quality content is also concentrated among those hubs. This implies that high-quality content has more competition and becomes obsolete more quickly compared to the case in which this content is uniformly distributed among authentic nodes. The same does not apply to content from bad actors because this content spreads uniformly among authentic users. (The scenario in which content from bad actors mostly concentrates among hubs is explored later.)
@@
Bad actors may maximize their message spread by combining various manipulation tactics. We systematically quantify the effects of these tactics through simulations varying the parameters for infiltration ($10^{-4}\leq\gamma\leq 10^{-1}$), deception ($0\leq\phi\leq 1$), and flooding ($1\leq\theta\leq 64$). See Methods for further details. Fig. 4(a,b,c) illustrates the effects of individual malicious tactics on the overall quality of information spreading through the network, compared to a baseline scenario without bad actors. We observe that infiltration is the most harmful manipulation tactic: when authentic agents have a $\gamma=10\%$ probability of following each bad actor, the average quality in the system is reduced to less than half. Flooding and deception have smaller effects. When low-quality content has $\theta=64$ times more exposure than authentic content, quality is reduced to less than 70%. When bad actors generate content with maximum appeal exclusively ($\phi=1$), quality is reduced to about 70%.
@@
Similarly, Fig. 4(d,e,f) shows the effects of pairs of tactics combined. Infiltration is dominant, but more harm can be done in combination with flooding or deception: the average quality is reduced to 40% when $\gamma=0.1$ and $\theta=64$ (Fig. 4(d)) or $\phi=1$ (Fig. 4(e)). Combining flooding and deception ($\theta=64,\phi=1$, Fig. 4(f)) only results in marginal loss of quality (below 70%). With all three tactics combined ($\gamma=0.1,\theta=64,\phi=1$, not shown), bad actors can further reduce the quality to 30%.
@@
Empirical evidence has shown that among fact-checked claims, low-quality content (debunked claims) tends to have larger retweet cascades than high-quality content (confirmed claims) (9, 45). SimSoM allows us to examine how different factors may contribute to such a virality pattern. Fig. 5 illustrates the effects of bad-actor tactics on the size of reshare cascades for content generated by both authentic agents (“high-quality”) and inauthentic agents (“low-quality”). We observe that low-quality cascade size is boosted by appeal. This is in line with the hypothesis that factors like novelty make false news more viral (9). However, low-quality cascades can be boosted most effectively through high bad-actor infiltration. Importantly, empirical analyses that do not reconstruct actual cascades from the data may underestimate intermediary amplification by inauthentic accounts, even when those accounts are removed as originators (cascade tree roots). On the other hand, our model lets us reconstruct the likely reshare cascades that include intermediary amplification by inauthentic accounts (see Methods).
Figure 5: Complementary cumulative distributions of reshare cascade sizes for low- and high-quality content, generated by inauthentic and authentic agents, respectively. The plots are based on 10 simulations. (a) Effect of bad-actor infiltration $\gamma$, with no flooding ($\theta=1$) or deception ($\phi=0$). (b) Effect of flooding $\theta$, with low infiltration ($\gamma=10^{-3}$) and no deception ($\phi=0$). (c) Effect of deception $\phi$, with low infiltration ($\gamma=10^{-3}$) and no flooding ($\theta=1$).
To date, empirical data from social media platforms has allowed researchers to measure reshare cascades, but not exposure. As a result, few studies focus on exposure (46). Using the SimSoM model, it is possible to reconstruct not only likely reshare networks but also exposure networks, thus estimating how many accounts are exposed to (i.e., view) a message even if they do not reshare it (see Methods). Fig. 6 compares the sizes of reshare and exposure cascades. In general, reshares underestimate exposures by roughly one order of magnitude. Excluding the smallest cascades and the largest ones (to avoid finite-size effects), we observe that exposure networks grow sub-linearly with reshare networks: $s_{v}\sim s_{r}^{\nu}$ where $s_{v}$ and $s_{r}$ are the exposure and reshare cascade sizes, respectively, and $\nu<1$ is the scaling exponent. This means that as messages go viral, exposures grow more slowly than reshares. The exponent is higher for low-quality ($\nu\approx 0.8$) than for high-quality content ($\nu\approx 0.6$), suggesting that for each extra reshare, messages posted by inauthentic agents gain more views.
@@
As an example, an adversary might target influentials based on the assumption that having such followers can multiply their impact — a message reshared by an influential account has a higher chance of going viral. Targeting influentials is well within the capability of bad actors and even automated accounts; the number of followers, often used as a proxy for influence (47), is public information on all social media platforms. A bad actor can easily interact with accounts having many followers by mentioning and/or following them (38, 48); other ploys include retweeting, quoting, and/or liking their tweets. There is empirical evidence of preferential targeting by bad actors that spread misinformation (1, 8). Targeting politically active accounts or habitual misinformation spreaders are also conceivable tactics.
An important question, then, is whether tactics targeting specific authentic accounts do in fact increase the manipulative power of bad actors. To explore this question, the model can be extended to account for various features of social media users. Here we consider five features that are available in the empirical data: number of followers (hubs tactic), propensity to share misinformation (misinformation tactic), political partisanship (partisanship tactic), or specific political leaning (liberal and conservative tactic). We then introduce a preferential targeting tactic for adversarial actors, according to which authentic agents have different probabilities of following inauthentic ones, proportional to one of these features. See Methods for details.
Figure 7: Effects of targeting tactics. (a) Average information quality resulting from each tactic, as well as the default random targeting, relative to the scenario without bad actors. We highlight significant differences calculated using The Mann–Whitney U test (**** for $p<10^{-4}$). (b) Suppression of quality in the empirical network when bad actors specifically target influential accounts (hubs), and when they target politically left- (liberal) and right-leaning (conservative) accounts. The network has $10^{3}$ authentic agents (purple nodes) and 50 inauthentic agents. Node size represents the number of followers. The darker an authentic agent node, the lower the quality of messages in their feed. Significant changes due to targeting tactics are only observed when bad-actor infiltration is sufficiently high, therefore we use $\gamma=10^{-1}$ in experiments for both panels.
@@
The harm of manipulation through the network also diminishes when bad actors target accounts sharing a lot of misinformation or with specific political leaning ($p<10^{-4}$). The echo-chamber structure of the empirical network (Fig. 7(b)) helps interpret the latter finding: low-quality messages get shared and become obsolete rapidly within one densely connected partisan cluster, sparing the rest of the network.